OpenStudy (anonymous):

Can you please help me with this? I don't get it. [9.02] What are the x-intercept(s) of the graph of y − x2 = 11x + 30?

OpenStudy (anonymous):

A couple of things first: One, it is good to get into the necessary habit of writing x^2, not x2, because x2 really means 2x. Two, best to re-write this as an equation: y = x^2 + 11x + 30 Done by subtracting x^2 from each side. Three, the x-intercepts are where y = 0, so you are dealing with: x^2 + 11x + 30 = 0 That can be factored. Are you able to do that?

OpenStudy (anonymous):

I'm not the best at it..

OpenStudy (anonymous):

I don't think I can. I'm not exactly sure how to do that.

OpenStudy (anonymous):

np. You are looking for 2 numbers that add up to 11 but when multiplied give 30. (x + 5)(x + 6) = 0

OpenStudy (anonymous):

So, your x-intercepts are x = -5 and x = -6. At those 2 values for x you have y = 0

OpenStudy (anonymous):

Okay. Thanks. :)

OpenStudy (anonymous):

You're quite welcome!

OpenStudy (anonymous):

Good luck in all of your studies.